# SoK: Membership Inference Attacks on LLMs are Rushing Nowhere (and How to Fix It)

### v2

@@ Abstract
Whether Large Language models (LLMs) memorize their training data and what this means, from privacy leakage to detecting copyright violations — has become a rapidly growing area of research over the last two years. In recent months, more than $10$ new methods have been proposed to perform Membership Inference Attacks (MIAs) against LLMs. Contrary to traditional MIAs which rely on fixed -but randomized- records or models, these methods are mostly evaluated on datasets collected post-hoc. Sets of members and non-members, used to evaluate the MIA, are constructed using informed guesses after the release of a model. This lack of randomization raises concerns of a distribution shift between members and non-members. In the first part of this paper, we review the literature on MIAs against LLMs. While most work focuses on sequence-level MIAs evaluated in post-hoc setups, we show that a range of target models, motivations and units of interest have been considered in the literature. We then quantify distribution shifts present in the $6$ datasets used in the literature, ranging from books to papers, using a bag of word classifier. Our analysis reveals that all of them suffer from severe distribution shifts. This challenges the validity of using such setups to measure LLM memorization and may undermine the benchmarking of recently proposed methods.
Yet, all hope might not be lost. In the second part of this work, we introduce important considerations to properly evaluate MIAs against LLMs and discuss potential ways forward: randomized test splits, injections of randomized (unique) sequences, randomized finetuning, and several post-hoc control methods. While each option comes with its advantages and limitations, we believe they collectively provide solid grounds to guide the development of MIA methods and study LLM memorization. We conclude by proposing and releasing comprehensive, easy-to-use benchmarks for sequence- and document-level MIAs against LLMs (see Table 6 and the datasets and code we release [^1] Link to the Github repository.. LLM memorization is an important and multi-faceted question, yet meaningful progress can only be achieved with the use of robust, independent benchmarks such as the ones we propose here.
1 Introduction
@@
The prevalence of Large Language Models (LLMs) [77, 3, 43] has sparked significant interest in studying LLM memorization to better understand how they acquire their capabilities and learn facts. LLMs can also be trained on confidential or sensitive personal data, making the study of LLM memorization critical from both an information security and a privacy perspective [9, 71, 15, 65, 16, 72]. Similar research helps assess whether LLMs are being trained on copyrighted material [86, 100, 70, 28] and the extent to which they might generate such content [46]. Further, to fairly evaluate newly released models, it is crucial to understand whether they are trained on (and memorize) evaluation benchmarks [66, 56, 78, 33, 22, 54, 25].
@@
Membership Inference Attacks (MIAs) have become a widely used tool to study the memorization of machine learning (ML) models [88, 12]. In an MIA, an attacker aims to infer whether a certain data sample has been used to train the target model. Recently, MIAs have also been proposed for LLMs [16, 106, 68, 86, 70, 109, 101, 62, 105, 108, 97, 73, 45], trying to infer whether a specific piece of text has been used to train an LLM. In most of this work, the unit of interest to infer membership has been a sequence (typically $\leq 256$ tokens), while other work has considered document [69, 86], to even dataset-level MIAs [67]. Beyond MIAs, prior work has also widely studied training data extraction from LLMs, either verbatim [13, 74, 112] or not [40, 80, 46], to measure LLM memorization.
@@
Independent of how MIAs against ML models (e.g. image classification) are developed, they are traditionally evaluated in randomized experiments [88, 12], instantiated as a privacy game between a model developer and an attacker. Ye et al. [104] distinguishes between privacy games for fixed records and for fixed models. For fixed records, multiple models are trained with the target randomly included or excluded (shadow modeling) [63, 64, 91, 41]. Fixed model scenarios use a single model trained on randomly sampled, IID records [76, 84, 98]. While feasible for traditional ML models, applying a proper privacy game for LLMs presents unique challenges. Indeed, the mere scale and cost of training prohibits the development of LLMs solely for the purpose of developing proper privacy games. Further, it remains uncertain how findings from small-scale MIA experiments translate to the scale of real-world LLMs.
@@
As a result, recent MIAs against LLMs have been evaluated in the absence of a randomized privacy game, relying instead on observational data collected post-hoc. Two concurrent works, published at ICLR [86] and USENIX Security [69], instantiate the membership inference task post-hoc; where an attacker aims to apply an MIA on a released LLM without access to the randomized training data distribution. Authors use publicly available datasets widely used for LLM training to collect likely members, and use data released on the same sources but after the LLM’s training data cutoff date to collect likely non-members. In this setup, both contributions report MIAs to work very well, reaching AUCs up to $0.8$. Further work has either used the exact same datasets, or similar datasets collected post-hoc, to develop and evaluate MIAs against LLMs [109, 101, 62, 105, 108, 97, 73, 45], all reaching high AUC values.
@@
However, recently, strong concerns have been raised on the lack of randomization in post-hoc collection of members and non-members. Duan et al. [27] first suggested that this setup introduces a distribution shift between members and non-members due to the temporal gap in data collection, a concern later echoed by Maini et al. [67]. This shift would lead membership classifiers to learn temporal differences in language rather than membership, potentially inflating AUC scores and leading to an incorrect understanding of LLM memorization.
Contribution. In this work, we provide an overview of the recent surge in research on MIAs against LLMs and quantify any distribution shift in the evaluation setups used. We find that the absence of randomization has both a substantial impact on the measured MIA performance and that this issue is widespread. We then carefully examine the options towards more rigorous MIA evaluation to, ultimately, improve our understanding of LLM memorization.
@@
First, we formalize the task of MIAs against LLMs, distinguishing between MIAs targeting LLM pretraining and finetuning, on the sequence-, document- and dataset-level. In Table 1, we review the literature, with a focus on the most recent advances (’23-’24) while including seminal initial work. MIA studies either merely evaluate MIAs from the literature on specific setups or also propose novel methods. We find that only in ’24, at least $10$ new MIA methods have been proposed. Next, we find that the primary motivation for such studies has evolved over time, with initial work mostly studying LLM finetuning for privacy and security reasons [15, 16, 71, 68, 65], while more recent work studies MIAs also for copyright and data contamination purposes, increasingly considering LLM pretraining. We then carefully examine which datasets have been considered to evaluate MIAs across the literature. We find that the majority of recent work does not consider randomized privacy games, and instead relies on data collected post-hoc (flagged with in Table 1).
@@
Second, across datasets collected post-hoc, we examine the distribution shift between members and non-members as flagged by prior work [27, 67] to understand its impact on previously reported results. We propose to train a bag of words classifier to distinguish members from non-members, without considering the model [^2] Concurrently with this work, Das et al. [21] applies ‘blind’ baselines to datasets used for MIA evaluation. Authors confirm that simple classifiers can effectively distinguish members from non-members without taking the target model into account.. We apply this to data collected post-hoc [69, 86, 57, 105] and confirm that the simple classifier yields an AUC often better than proposed MIAs. This shows that there exists a serious distribution shift in language between the collected members and non-members - making it impossible to exclusively attribute the MIA performance to LLM memorization. This questions the effectiveness of newly proposed MIAs [109, 101, 62, 105, 108, 97, 73, 45] (latest one from September, 2024), potentially leading to false conclusions on LLM memorization.
@@
Third, we discuss what a randomized setup to evaluate MIAs against LLMs entails. The cost of LLM development prohibits training multiple models for MIA evaluation, leaving only the privacy game for a fixed model as a practical option. Before examining the potential solutions, we elaborate on two core challenges. First, the inherent overlap between texts collected from the internet challenges the binary notion of membership. Prior work found significant overlap between randomized splits from the training data, either trying to achieve a clean model validation set [10, 99, 26, 95] or a useful set of members and non-members to evaluate MIAs [27]. Second, we elaborate on the unit of interest considered in MIAs against LLMs, ranging from a sequence to an entire dataset.
@@
Lastly, we carefully examine the approaches available for rigorous evaluation of MIAs against LLMs in the future. We consider the following $4$ options and elaborate on their advantages and limitations:
@@
    The use of the randomized split of train and test data released for open-source LLMs as members and non-members for MIA evaluation, as suggested by prior work [55, 27, 67].
@@
    The application of control methods to make evaluation and calibration of MIAs against LLMs feasible on data collected post-hoc, either with a skewed prior [42, 47], de-biasing techniques [29] or techniques inspired by Regression Discontinuity Design (RDD) [50].
We conclude with an overview of recommended approaches to benchmark newly proposed MIAs in the future (Table 6). We hope this overview contributes to a better understanding of the recent advances in MIAs against LLMs and provides a foundation for future research to rigorously evaluate and interpret MIA results.
@@
    While MIMIR originates from a randomized setup, subsquent deduplication [27] could introduce a distribution shift (Sec. 6.1).
@@
    SHA-hashes extracted from the StackExchange datasets, likely doesn’t suffer from the distribution shift.
Table 1: Overview of prior work studying MIAs against LLMs. We distinguish target models, tasks, motivations and datasets.
2 Preliminary
Threat model. We here consider a target LLM $\mathcal{M}$ trained on dataset $\mathcal{D}$, with tokenizer $T$. The goal of the attacker $\mathcal{A}$ is now to leverage access to the LLM $\mathcal{M}$ to infer whether a specific piece of text has been part of its training data $\mathcal{D}$.
We note that MIAs against LLMs have been proposed across varying assumptions made for the attacker $\mathcal{A}$, from access to the model weights [55, 62], to only query access with model logits [106, 16, 68, 86, 69, 109, 101, 105, 108, 97, 73] or merely generated text [28, 45]. Some work also assumes access to additional non-member data from the same distribution as the target sample [101, 62, 97]. In this work, we focus on the evaluation of MIAs against LLMs, independent of the specific threat model applied.
Dataset Structure for LLM Training. We formalize the LLM training dataset $\mathcal{D}$ as a collection of $N$ documents: $\mathcal{D}=\{d_{1},d_{2},\ldots,d_{N}\}$, where each document $d_{i}$ is composed of a sequence of tokens.
A document is an ordered sequence of tokens, $d_{i}=(t_{1},t_{2},...,t_{N_{d_{i}}})$, with $N_{d_{i}}$ the length of document $d_{i}$. We consider documents originating from a single, identifiable source which are coherent and semantically meaningful when considered in isolation. Examples include a news article, blogpost, book, academic paper, piece of code, etc. The document length can vary widely. For a single book this can be up to more than 100k tokens, while for social media posts this can be under $100$ tokens. The document length is a property of the document itself and is independent of whether the entire document fits into the LLM’s context window.
@@ 2 Preliminary
LLM (pre)training. Tokenizer $T$ maps a piece of text $S_{ij}$ to a sequence of tokens $T(S_{ij})=s_{ij}=(t_{a},t_{a+1},\ldots,t_{b})$, where each token belongs to a vocabulary $\mathcal{V}$ of size $|\mathcal{V}|=V$.
@@
Given a sequence of tokens (referred to as the context), auto-regressive language models predict the probability distribution for the next token over all tokens in vocabulary $\mathcal{V}$. This capability is acquired during training on the entire dataset $\mathcal{D}$, where the model minimizes the loss, generally measured using the cross-entropy function, for next-token prediction. For a model $\mathcal{M}$ with parameters $\theta$, we define the loss computed on sequence $s_{ij}$ as $\mathcal{L}_{\mathcal{M}}(s_{ij})=-\frac{1}{b-a+1}\sum_{k=a}^{b}\log\left(\mathcal{M}_{\theta}(t_{k}|t_{a},\ldots,t_{k-1})\right)$, where $\mathcal{M}_{\theta}(t_{k}|t_{a},\ldots,t_{k-1})$ represents the probability predicted by the model for the true token $t_{k}$ given the preceding tokens.
@@
In this work, we exclusively consider models trained on causal language modeling objectives. This includes both base model versions — models trained on large quantities of unsupervised data — and versions that have been further fine-tuned on supervised task-specific data to unlock further abilities (instruction following, chat, specific tasks). In Table 1, we refer to these target models as pretrained and finetuned, respectively. Models trained with more complex alignment techniques such as reinforcement learning methods fall outside the scope of this work.
@@
Sequence-level MIAs. The unit of interest in MIAs against LLMs has most commonly been a sequence. A sequence-level MIA thus aims to determine whether a given sequence $s_{t}$ has been used to train target LLM $\mathcal{M}$. Given an LLM $\mathcal{M}$ trained on dataset $\mathcal{D}$ and a target sequence $s_{t}$, a sequence-level MIA is a scoring function $f_{\text{seq}}$ such that $f_{\text{seq}}(\mathcal{M},s_{t})$ outputs a continuous value representing the likelihood that $s_{t}$ was included in $\mathcal{D}$. A higher score indicates a higher likelihood of membership. To make a binary prediction $\hat{y}$, a threshold $\tau$ can be applied $\hat{y}_{\text{seq}}(\mathcal{M},s_{t})=1\text{ if }f_{\text{seq}}(\mathcal{M},s_{t})\geq\tau_{\text{seq}}\text{, and }0\text{ otherwise}$. The MIA performance can then be determined with the true label for membership $y$, $y_{\text{seq}}(\mathcal{M},s_{t})=1\text{ if }\exists d_{i}\in\mathcal{D},s_{t}\in d_{i}\text{, and }0\text{ otherwise}$.
MIAs are typically evaluated on the scoring function $f_{\text{seq}}$ using classification metrics that do not rely on a particular threshold, such as area under the receiver operating characteristic curve (AUC) or true positive rate (TPR) at low false positive rate (FPR) [12].
@@
Most prior studies on MIAs against LLMs have considered sequences of a fixed, often truncated length significantly shorter than most documents. Carlini et al. [16] use sequences of $256$ tokens, while Shi et al. [86] and subsequent works [109, 101, 62, 105, 108, 97, 73, 45] consider random excerpts between $32$ and $256$ tokens and Meeus et al. [70] considers up to $100$ synthetically generated tokens. We thus primarily focus on sequence-level MIAs and discuss other units of interest in Sec. 5.2 and 6.
@@
An MIA against an LLM thus comes down to designing the scoring function $f_{\text{seq}}(\mathcal{M},s_{t})$ which assesses the likelihood that a sequence $s_{t}$ was observed by model $\mathcal{M}$. Intuitively, model loss can be used as a naive scoring function (i.e. $f_{\text{seq}}(\mathcal{M},s_{t})=-\mathcal{L}_{\mathcal{M}}(s_{t})$[106], as model training minimizes entropy values for sequences included in the training data. Notably, individual sequences are not seen in isolation during training, as they are likely conditioned on preceding context. Especially for sequences smaller than documents (or data grouped during training), this might impact the effectiveness of using the model loss as a scoring function in sequence-level MIAs against LLMs compared to other ML models. Further MIAs against LLMs have proposed alternative scoring functions, for instance, by normalizing the model loss with respect to how ’normal’ a piece of text is [16], by contrasting the model loss with the loss of neighboring samples [68] or by only focusing on the lowest predicted probabilities [86].
3 Literature review

### v3

# Anonymous authors    Matthieu Meeus1, Igor Shilov1, Shubham Jain2, Manuel Faysse3, Marek Rei1, Yves-Alexandre de Montjoye1 1Imperial College London 2Sense Street

@@ Abstract
Whether Large Language models (LLMs) memorize their training data and what this means, from measuring privacy leakage to detecting copyright violations, has become a rapidly growing area of research. In the last few months, more than 10 new methods have been proposed to perform Membership Inference Attacks (MIAs) against LLMs. Contrary to traditional MIAs which rely on fixed—but randomized—records or models, these methods are mostly trained and tested on datasets collected post-hoc. Sets of members and non-members, used to evaluate the MIA, are constructed using informed guesses after the release of a model. This lack of randomization raises concerns of a distribution shift between members and non-members. In this work, we first extensively review the literature on MIAs against LLMs and show that, while most work focuses on sequence-level MIAs evaluated in post-hoc setups, a range of target models, motivations and units of interest are considered. We then quantify distribution shifts present in 6 datasets used in the literature, ranging from books to papers using a model-less bag of word classifier and compare the model-less results to the MIA. Our analysis shows that all of these datasets constructed post-hoc suffer from strong distribution shifts. These shifts invalidate the claims of LLMs memorizing strongly in real-world scenarios and, potentially, also the methodological contributions of the recent papers based on these datasets. Yet, all hope might not be lost. In the second part of this work, we introduce important considerations to properly evaluate MIAs against LLMs and discuss, in turn, potential ways forwards: randomized test splits, injections of randomized (unique) sequences, randomized fine-tuning, and several post-hoc control methods. While each option comes with its advantages and limitations, we believe they collectively provide solid grounds to guide the development of MIA methods and study LLM memorization. We conclude with an overview of recommended approaches to benchmark sequence-level and document-level MIAs against LLMs.
I Introduction
@@
The prevalence of Large Language Models (LLMs) [76, 2, 43] has sparked significant interest in studying LLM memorization to better understand how they acquire their capabilities and learn facts. LLMs can also be trained on confidential or sensitive personal data, making the study of LLM memorization critical from both an information security and a privacy perspective [8, 70, 14, 63, 15, 71]. Similarly, research helps assess whether LLMs are being trained on copyrighted material [85, 99, 69, 28] and the extent to which they might generate such content [46]. Further, to fairly evaluate newly released models, it is crucial to understand whether they are trained on (and memorize) evaluation benchmarks [64, 55, 77, 33, 22, 53, 25].
@@
Membership Inference Attacks (MIAs) have become a widely used tool to study the memorization of machine learning (ML) models [87, 11]. In an MIA, an attacker aims to infer whether a certain data sample has been used to train the target model. Recently, MIAs have also been proposed for LLMs [15, 105, 66, 85, 69, 109, 100, 60, 104, 107, 96, 72, 45] to infer whether a specific piece of text has been used to train an LLM. While most of these works have focused on sequences as the unit of interest to infer membership (typically $\leq 256$ tokens), other works have considered document [68, 85], and dataset-level MIAs [65]. Beyond MIAs, prior work has also widely studied training data extraction from LLMs, either verbatim [12, 73, 112] or not [40, 79, 46], to measure LLM memorization.
@@
Independently of how MIAs against ML models (e.g. image classification) are developed, they are traditionally evaluated in randomized experiments [87, 11], instantiated as a privacy game between a model developer and an attacker. Ye et al. [103] proposes to distinguish between privacy games with fixed records and fixed models. For fixed records, multiple models are trained with the target randomly included or excluded [61, 62, 90, 41]. Fixed model scenarios use a single model trained on randomly sampled, IID records [75, 83, 97]. While feasible for traditional ML models, evaluating MIAs using a privacy game for LLMs presents unique challenges. On the one hand, the mere scale and cost of training indeed prohibits the development of LLMs solely for the purpose of playing a privacy game. On the other hand, it remains uncertain how findings from small-scale MIA experiments translate to the scale of real-world LLMs.
@@
As a result, recent MIAs against LLMs have been evaluated using observational data collected post-hoc. Two concurrent works, published at ICLR [85] and USENIX Security [68] in 2024, first instantiate the membership inference task post-hoc; where an attacker aims to apply an MIA on a released LLM without access to the randomized training data distribution. Authors both use publicly available datasets widely used for LLM training to collect likely members, and use data released on the same sources but after the LLM’s training data cutoff date to collect likely non-members. Both contributions report MIAs to work very well, with AUCs up to $0.8$. Since then, numerous pieces of work used either the same datasets of members and non-members, or similar datasets collected post-hoc, to develop and evaluate MIAs against LLMs [109, 100, 60, 104, 107, 96, 72, 45], all reporting high AUCs.
@@
In parallel with new MIAs being developed, strong concerns have been raised on the lack of randomization in post-hoc collection of members and non-members. Duan et al. [27] first suggested that this setup introduces a distribution shift between members and non-members due to the temporal gap in data collection. These concerns were shared by Maini et al. [65] and later also by Das et al. [21] [^1] Das et al. [21] is concurrent with this work (for a comparison see Sec. VII).. Such distribution shifts would lead membership classifiers to learn temporal differences in language rather than membership, potentially inflating AUC scores and leading to an incorrect understanding of LLM memorization.
In this work, we provide an overview of the recent surge in research on MIAs against LLMs, identifying numerous works using datasets collected post-hoc used for MIA evaluation. With a bag of words classifier, we then identify serious distribution shifts present in these datasets. Lastly, we carefully examine the requirements for rigorous MIA evaluation and solutions moving forward.
@@
First, we formalize the task of MIAs against LLMs, distinguishing between MIAs targeting LLM pretraining and finetuning, on the sequence-, document- and dataset-level. In Table I, we review the literature, with a focus on the most recent advances (’23-’24) while including seminal initial work. We find that only in ’24, at least $10$ new MIA methods have been proposed. Next, we find that the primary motivation for such studies has evolved over time, with initial work mostly studying LLM finetuning for privacy and security reasons [14, 15, 70, 66, 63], while more recent work studies MIAs to identify copyright violations or data contamination, increasingly considering LLM pretraining. We then carefully examine which datasets have been considered to evaluate MIAs across the literature. We find that the majority of recent work does not consider randomized privacy games, and instead relies on data collected post-hoc (flagged with in Table I).
@@
Second, we introduce a bag of words classifier to examine the distribution shift between members and non-members as flagged by prior [27, 65] and concurrent work [21]. We train the bag of words classifier to distinguish members from non-members, without considering the target model and use its performance to quantify the distribution shift and understand its impact on previously reported results. We apply this to $6$ datasets collected post-hoc [68, 85, 56, 104] (used in over $10$ papers for MIA evaluation) and confirm that the simple classifier yields an AUC often better than proposed MIAs. This shows that there exists such a serious distribution shift in language between the collected members and non-members that it is impossible to exclusively attribute the MIA performance to LLM memorization. This questions the effectiveness of newly proposed MIAs [109, 100, 60, 104, 107, 96, 72, 45] (latest one from September, 2024), potentially leading to false conclusions on LLM memorization.
@@
Third, before examining the potential solutions to evaluate MIAs against LLMs, we elaborate on two core challenges. First, the inherent overlap between texts collected from the internet challenges the binary notion of membership. Prior work found significant overlap between randomized splits from the training data, either trying to achieve a clean model validation set [9, 98, 26, 94] or a useful set of members and non-members to evaluate MIAs [27]. Second, we elaborate on the unit of interest considered in MIAs against LLMs, ranging from a sequence to an entire dataset.
@@
Lastly, we examine the approaches available for rigorous evaluation of MIAs against LLMs in the future. We consider the following $4$ options and elaborate on their advantages and limitations:
@@
    The use of the randomized split of train and test data released for open-source LLMs as members and non-members, as suggested by prior work [54, 27, 65].
@@
    The application of control methods to make evaluation and calibration of MIAs against LLMs feasible on data collected post-hoc, either with a skewed prior [42, 47] or de-biasing techniques [29] as proposed by recent work, or as we here propose using techniques inspired by Regression Discontinuity Design (RDD) [50].
We conclude with an overview of recommended approaches to benchmark newly proposed sequence and document-level MIAs in the future (Table VI). We hope this overview contributes to a better understanding of the recent advances in MIAs against LLMs and provides a solid foundation for future research to rigorously evaluate and interpret MIA results.
TABLE I: Overview of prior work studying MIAs against LLMs. We distinguish target models, tasks, motivations and datasets.
@@
    While MIMIR originates from a randomized setup, subsquent deduplication [27] could introduce a distribution shift (Sec. VI-A).
@@
    SHA-hashes extracted from the StackExchange datasets, likely does not suffer from the distribution shift.
II Preliminary
Threat model. We consider a target LLM $\mathcal{M}$ trained on dataset $\mathcal{D}$, with tokenizer $T$. The goal of the attacker $\mathcal{A}$ is to leverage access to the LLM $\mathcal{M}$ to infer whether a specific piece of text has been part of the training data $\mathcal{D}$.
MIAs against LLMs have been proposed across varying assumptions made for the attacker $\mathcal{A}$, from access to model weights (white-box) [54, 60], to only query access with model logits (black-box) [105, 15, 66, 85, 68, 109, 100, 104, 107, 96, 72] or merely generated text [28, 45] (no-box). Some work also assumes access to additional non-member data from the same distribution as the target sample [100, 60, 96]. We here focus on the evaluation of MIAs against LLMs, independent of the threat model used.
Dataset Structure for LLM Training. We formalize the LLM training dataset $\mathcal{D}$ as a collection of $N$ documents: $\mathcal{D}=\{d_{1},d_{2},\ldots,d_{N}\}$, where each document $d_{i}$ is composed of a sequence of tokens. A document is an ordered sequence of tokens, $d_{i}=(t_{1},t_{2},...,t_{N_{d_{i}}})$, with $N_{d_{i}}$ the length of document $d_{i}$. We consider documents originating from a single, identifiable source which are coherent and semantically meaningful when considered in isolation. Examples include a news article, blogpost, book, academic paper or piece of code. The document length can vary widely and is a property of the document itself, thus independent of whether the entire document fits into the LLM’s context window. For a single book this can be up to more than 100k tokens, while for social media posts this can be under $100$ tokens.
@@ II Preliminary
LLM training. Tokenizer $T$ maps a piece of text $S_{ij}$ to a sequence of tokens $T(S_{ij})=s_{ij}=(t_{a},t_{a+1},\ldots,t_{b})$, where each token belongs to a vocabulary $\mathcal{V}$ of size $|\mathcal{V}|=V$. Given a sequence of tokens (referred to as the context), auto-regressive language models predict the probability distribution for the next token over all tokens in vocabulary $\mathcal{V}$. This capability is acquired during training on the entire dataset $\mathcal{D}$, where the model minimizes the loss, generally cross-entropy loss, for next-token prediction. For a model $\mathcal{M}$ with parameters $\theta$, we define the loss computed on sequence $s_{ij}$ as $\mathcal{L}_{\mathcal{M}}(s_{ij})=-\frac{1}{b-a+1}\sum_{k=a}^{b}\log\left(\mathcal{M}_{\theta}(t_{k}|t_{a},\ldots,t_{k-1})\right)$, where $\mathcal{M}_{\theta}(t_{k}|t_{a},\ldots,t_{k-1})$ represents the probability predicted by the model for the true token $t_{k}$ given the preceding tokens.
@@
In this work, we exclusively consider models trained on causal language modeling objectives. This includes both base model versions—trained on large quantities of unsupervised data—and versions that have been further fine-tuned on supervised task-specific data to unlock further abilities (instruction following, chat, specific tasks). In Table I, we refer to these models as pretrained and finetuned, respectively. In this work, we do not consider models trained with more complex alignment techniques such as reinforcement learning methods.
@@
Sequence-level MIAs. The unit of interest in MIAs against LLMs has most commonly been a sequence. A sequence-level MIA aims to determine whether a given sequence $s_{t}$ has been used to train target LLM $\mathcal{M}$. Given an LLM $\mathcal{M}$ trained on dataset $\mathcal{D}$ and a target sequence $s_{t}$, a sequence-level MIA is a scoring function $f_{\text{seq}}$ such that $f_{\text{seq}}(\mathcal{M},s_{t})$ outputs a continuous value representing the likelihood that $s_{t}$ was included in $\mathcal{D}$. A higher score indicates a higher likelihood of membership. To make a binary prediction $\hat{y}$, a threshold $\tau$ can be applied $\hat{y}_{\text{seq}}(\mathcal{M},s_{t})=1\text{ if }f_{\text{seq}}(\mathcal{M},s_{t})\geq\tau_{\text{seq}}\text{, and }0\text{ otherwise}$. The MIA performance is then determined with the true label for membership $y$, $y_{\text{seq}}(\mathcal{M},s_{t})=1\text{ if }\exists d_{i}\in\mathcal{D},s_{t}\in d_{i}\text{, and }0\text{ otherwise}$. MIAs are typically evaluated on the scoring function $f_{\text{seq}}$ using classification metrics that do not rely on a particular threshold, such as area under the receiver operating characteristic curve (AUC) or true positive rate (TPR) at low false positive rate (FPR) [11].
@@
Most prior studies on MIAs against LLMs have considered sequences of a fixed, often truncated length significantly shorter than most documents. Carlini et al. [15] use sequences of $256$ tokens, while Shi et al. [85] and subsequent works [109, 100, 60, 104, 107, 96, 72, 45] consider random excerpts between $32$ and $256$ tokens and Meeus et al. [69] considers up to $100$ synthetically generated tokens. We thus primarily focus on sequence-level MIAs and discuss other units of interest (i.e. documents and datasets, typically longer than sequences) in Sec. V-B and VI.
@@
An MIA against an LLM thus involves designing the scoring function $f_{\text{seq}}(\mathcal{M},s_{t})$ which assesses the likelihood that a sequence $s_{t}$ was observed by model $\mathcal{M}$. For instance, model loss can be used as a naive scoring function (i.e. $f_{\text{seq}}(\mathcal{M},s_{t})=-\mathcal{L}_{\mathcal{M}}(s_{t})$[105], as model training minimizes entropy values for sequences included in the training data. Individual sequences are, however, not seen in isolation during training, as they are likely conditioned on preceding context. Especially for sequences smaller than documents (or data grouped during training), this might impact the effectiveness of using the model loss as a scoring function in sequence-level MIAs against LLMs compared to other ML models. MIAs against LLMs have thus proposed more advanced scoring functions, by, for instance, normalizing the model loss with respect to how ’normal’ a piece of text is [15], contrasting the model loss with the loss of neighboring samples [66] or by only focusing on the lowest predicted probabilities [85].
III Literature review